Given the following information: utility function is U(x,y)=〖4x〗^0.5 y^0.5 , price of good X is N$5, the price of good Y is N$10 and the consumer income N$400.
What is the level of quantity demanded of good X when the price of good X and Y is N$4 and N$10 respectively? Let us assume good X is on the x-axis.
1
Expert's answer
2020-06-29T14:44:27-0400

We have a utility function:



"E(x,y) = 4x^{0.5}y^{0.5}"

A consumer with this utility function will maximize her utility by producing at the point where:



"\dfrac{MU_x}{MU_y} = \dfrac{P_x}{P_y}"

From the utility function:



"MU_x = \dfrac{\delta U(x,y)}{\delta x} = 2x^{-0.5}y^{0.5}\\[0.3cm]
MU_y = \dfrac{\delta U(x,y)}{\delta y} = 2x^{0.5}y^{-0.5}"

The price of good x is $5 and the price of good y is $10. Therefore:



"\dfrac{2x^{0.5}y^{-0.5}}{2x^{-0.5}y^{0.5}} = \dfrac{5}{10}\\[0.3cm]
\dfrac{y}{x} = \dfrac{1}{2}\\[0.3cm]
x = 2y............(i)"

The consumer's income is $400. Thus, the budget line is:



"400 = 5x + 10y"

Substituting equation (i) into the budget constraint, we get:


"400 = 5(2y) + 10y\\[0.3cm]
400 = 20y\\[0.3cm]
\color{red}{y^* = 20}"

Since "x = 2y" , then:



"x^* = 2(20)\\[0.3cm]
\color{red}{x^* = 40}"